# Adaptive Accelerated Composite Minimization

### v1

@@ 5. Conclusion and Future Directions
This research proposes a new stepsize rule for non-accelerated and accelerated first-order methods for convex composite optimization. The proposed approach provably achieves the convergence optimal rate of $\mathcal{O}(k^{-2})$ for accelerated methods. Compared to other stepsize rules, our proposed scheme is implicit but only requires zeroth-order information to compute a suitable stepsize. Possible extension to this work could be the $min\,max$ problem [26, 41, 10], stochastic gradient descent optimization [21], variational inequalities [27, 19, 25], and non-convex optimization [18, 39, 1]. To mitigate the fact that the stepsize in Alg. 2 is non-increasing, one could re-initialize it finitely many times after a predefined number of iterations. This technique is similar to the restarted accelerated gradient method [35, 20, 40] and it is left as future work.
The authors of a recent work [22] demonstrate that iterations generated by Forward–Backward splitting methods, which include several variants (e.g., ISTA, FISTA), lie on active manifolds within a finite number of iterations and enter a regime of local linear convergence. However, their analysis are based on the standard assumption that the stepsize belongs to $[0,2/L]$. Departing from this convention is one of the future areas of interest. The authors in [22] also establish and explain why FISTA locally oscillates and can be slower than ISTA under the same standard assumption $\lambda\in[0,2/L]$. We observe similar behavior in our simulation results (Figures 3a, 5b, 6). Analyzing this behavior for $\lambda>2/L$, which can also occur in our proposed stepsize rules, could be another future direction.

### v2

@@ 5. Conclusion and Future Directions
This research proposes a new stepsize rule for non-accelerated and accelerated first-order methods for convex composite optimization. The proposed approach provably achieves the convergence optimal rate of $\mathcal{O}(k^{-2})$ for accelerated methods. Compared to other stepsize rules, our proposed scheme is implicit but only requires zeroth-order information to compute a suitable stepsize. Possible extension to this work could be the minmaxmin\,max problem [mokhtari2020convergence, yoon2021accelerated, boct2022accelerated], stochastic gradient descent optimization [lacoste2012simpler], variational inequalities [nemirovski2004prox, juditsky2011solving, malitsky2020forward], and non-convex optimization [jin2018accelerated, vladarean2021first, alacaoglu2021convergence]. To mitigate the fact that the stepsize in Alg. 2 is non-increasing, one could re-initialize it finitely many times after a predefined number of iterations. This technique is similar to the restarted accelerated gradient method [Emmanuel2015, Kim2018, yang2018rsg] and it is left as future work.
The authors of a recent work [liang2017activity] demonstrate that iterations generated by Forward–Backward splitting methods, which include several variants (e.g., ISTA, FISTA), lie on active manifolds within a finite number of iterations and enter a regime of local linear convergence. However, their analysis are based on the standard assumption that the stepsize belongs to $[0,2/L]$. Departing from this convention is one of the future areas of interest. The authors in [liang2017activity] also establish and explain why FISTA locally oscillates and can be slower than ISTA under the same standard assumption $\lambda\in[0,2/L]$. We observe similar behavior in our simulation results (Figures 3a, 5b, 6). Analyzing this behavior for $\lambda>2/L$, which can also occur in our proposed stepsize rules, could be another future direction.